The length of one diagonal of a rhombus is 10in and the area is 70in find the length of the other diagonal
Answer by jim_thompson5910(35256)

A = p*q/2

70 = 10*q/2

70*2 = 10q

140 = 10q

140/10 = q

14 = q

q = 14

So the other diagonal is 14 inches.